# Non-uniqueness of the steady state for run-and-tumble particles with a double-well interaction potential

### v1

@@ Abstract
We study $N$ run-and-tumble particles (RTPs) in one dimension interacting via a double-well potential $W(r)=-k_{0}\,r^{2}/2+g\,r^{4}/4$, which is repulsive at short interparticle distance $r$ and attractive at large distance. At large time, the system forms a bound state where the density of particles has a finite support. We focus on the determination of the total density of particles in the stationary state $\rho_{s}(x)$, in the limit $N\to+\infty$. We obtain an explicit expression for $\rho_{s}(x)$ as a function of the ”renormalized” interaction parameter $k=k_{0}-3m_{2}$ where $m_{2}$ is the second moment of $\rho_{s}(x)$. Interestingly, this stationary solution exhibits a transition between a connected and a disconnected support for a certain value of $k$, which has no equivalent in the case of Brownian particles. Analyzing in detail the expression of the stationary density in the two cases, we find a variety of regimes characterized by different behaviors near the edges of the support and around $x=0$. Furthermore, by studying the relation between $k$ and $k_{0}$, we find that the mapping $k_{0}\to k$ becomes multi-valued below a certain value of the tumbling rate $\gamma$ of the RTPs for some range of values of $k_{0}$ near the transition, implying the existence of two stable solutions. Finally, we show that in the case of a disconnected support, it is possible to observe steady states where the density $\rho_{s}(x)$ is not symmetric, characterized by a third moment $m_{3}$ which can take a continuous range of values. All our analytical predictions are in good agreement with numerical simulations already for systems of $N=100$ particles. The non-uniqueness of the stationary state is a particular feature of this model in the presence of active (RTP) noise, which contrasts with the uniqueness of the Gibbs equilibrium for Brownian particles. We argue that these results are also relevant for a class of more realistic interactions with both an attractive and a repulsive part, but which decay at infinity.
@@ I.1 Overview
In this paper we consider one-dimensional RTPs interacting via a pairwise polynomial potential of the form $W(r)=-k_{0}\,r^{2}/2+g\,r^{4}/4$, which for $k_{0},g>0$ is repulsive at short distance and attractive at large distance. Although the interaction force $W^{\prime}(r)$ diverges at large distance, which is a priori an unrealistic feature (see however below), it has two advantages from a computational perspective. First, $W^{\prime}(0)=0$, hence it allows for particle crossings and therefore allows to use the extension of the Dean-Kawasaki method Dean ; Kawa to the RTP (while it was shown that this method fails to describe single-file active particles TouzoDBM2023 ). Second, it is simple enough so that the self-consistent equation obeyed by the stationary density in the large $N$ limit, derived in activeRDshort , can be solved analytically. We find that in the steady state the system forms a bound state with a bounded support, which exhibits a transition between a joint support and a disconnected support where particles spontaneously split into two groups. Furthermore, remarkably, we find that the steady state is not unique, and that in some range of parameters it exhibits bistability, as well as a breaking of the parity symmetry. The steady state which is reached at large time is found to depend on some features of the initial conditions. By contrast, both of these features are absent in the case of Brownian particles, for which the Gibbs equilibrium is unique.
@@ I.3 Main results
In order to test these analytical predictions, obtained in the limit $N\to+\infty$, we have performed numerical simulations of the equation of motion (1). The stationary density is determined by averaging over a large time window. We find a good agreement between theory and numerics already for $N=100$ (see Fig. 3). In particular, both the bistability (see Fig. 5) and the existence of asymmetric steady-states (see Fig. 8) are confirmed by these numerical results. The case of small values of $N$ is discussed in Appendix B.
@@ II.3.1 Relation between $k$ and $k_{0}$
Figure 4: Top left: Plot of $m_{2}$ versus $k$ for different values of $\gamma$. In the limit $\gamma\to 0^{+}$ (black curve), $m_{2}(k)$ has a simple expression given in (II.3.3). In this case, the function $m_{2}(k)$ is discontinuous at $k_{c}$, but it is smooth for any $\gamma>0$. Top right: Same plot zoomed around $k_{c}$, and for values of $\gamma$ close to $\gamma_{c}$. While there seems to be a cusp at $k_{c}$ for small $\gamma>0$ when looking at the large scales, the function $m_{2}(k)$ appears to be smooth when zooming sufficiently. Bottom left: Plot of $k$ versus $k_{0}$. When $\gamma$ is smaller than some critical value $\gamma_{c}=0.1787369$..., the function $k_{0}(k)$ becomes non-monotonous close to $k_{c}$. The inverse function $k(k_{0})$ thus becomes multi-valued, leading to the coexistence of 2 stable and one unstable steady states for the same value of $k_{0}$, corresponding to different values of $k$. The value $\gamma_{c}=0.1787369$... was computed numerically, using as a criterion that the minimum of $\frac{dk_{0}}{dk}$ vs $k$ vanishes at $\gamma_{c}$ (it is negative for $\gamma<\gamma_{c}$ and positive for $\gamma>\gamma_{c}$). For $\gamma=0^{+}$, the bistability occurs in the interval $k_{0}\in[6.614585...,9.449407$...]. Bottom right: Same plot zoomed around $k_{c}$, and for values of $\gamma$ close to $\gamma_{c}$. The region where the function $k_{0}(k)$ is non-monotonous seems to be entirely located at $k<k_{c}$.
@@ II.3.3 Bistability
$m_{2}(k)=\frac{1}{2}(y_{1}(k)^{2}+y_{3}(k)^{2})\quad,\quad k\geq k_{c}\;,$
@@
As a side remark, we note that in the limit $\gamma\to 0^{+}$, the values of $y_{1}$ and $y_{3}$ can be obtained explicitly as a function of the bare interaction parameter $k_{0}$ by inserting the relation (II.3.3) into the cubic equation (25) (using that $k=k_{0}-3m_{2}$). For $k<k_{c}$ this gives that $y_{1}$ is the only real root of
@@ III.4 Numerical results
We have tested some of the predictions above in numerical simulations (for $N=100$ particles). One finds that indeed starting from an asymmetric initial condition, e.g. separating the particles into two groups of unequal size at time $t=0$, one can obtain a non-symmetric stationary density, which is well described by the expression (78)-(79). See Fig. 8 for an example. In addition we have studied numerically small values of $N$ in the Appendix B and also found some asymmetry for $N=5$.
@@ IV Conclusion
There are several open problems even for this simple type of model. For instance one would like to be able to solve for the dynamics and obtain the basin of attraction of each steady state. Another direction is to extend this to higher order degree polynomials (or more general functions) and determine the co-dimension of the stationary manifold.
@@
Acknowledgments. We acknowledge support from ANR Grant No. ANR- 23-CE30-0020-01 EDIPS.
@@ Appendix A Self-consistent equation for the stationary density
$\partial_{t}\rho_{s}=\partial_{x}[(-v_{0}\rho_{d}-\tilde{F}(x,t)\rho_{s}]\;,$
@@
The equations (A) are identical to the equations of a single RTP, $N=1$, in a effective time dependent force field $\tilde{F}(x,t)$ which depends itself on the time dependent density. One can thus use the known results for this simpler problem, whenever they are available.

### v2

@@ Abstract
We study $N$ run-and-tumble particles (RTPs) in one dimension interacting via a double-well pairwise potential $W(r)=-k_{0}\,r^{2}/2+g\,r^{4}/4$, which is repulsive at short interparticle distance $r$ and attractive at large distance. At large time, the system forms a bound state where the density of particles has a finite support. We focus on the determination of the total density of particles in the stationary state $\rho_{s}(x)$, in the limit $N\to+\infty$. We obtain an explicit expression for $\rho_{s}(x)$ as a function of the “renormalized” interaction parameter $k=k_{0}-3m_{2}$ where $m_{2}$ is the second moment of $\rho_{s}(x)$. Interestingly, this stationary solution exhibits a transition between a connected and a disconnected support for a certain value of $k$, which has no equivalent in the case of Brownian particles. Analyzing in detail the expression of the stationary density in the two cases, we find a variety of regimes characterized by different behaviors near the edges of the support and around $x=0$. Furthermore, by studying the relation between $k$ and $k_{0}$, we find that the mapping $k_{0}\to k$ becomes multi-valued below a certain value of the tumbling rate $\gamma$ of the RTPs for some range of values of $k_{0}$ near the transition, implying the existence of two stable solutions. Finally, we show that in the case of a disconnected support, it is possible to observe steady states where the density $\rho_{s}(x)$ is not symmetric, characterized by a third moment $m_{3}$ which can take a continuous range of values. All our analytical predictions are in good agreement with numerical simulations already for systems of $N=100$ particles. The non-uniqueness of the stationary state is a particular feature of this model in the presence of active (RTP) noise, which contrasts with the uniqueness of the Gibbs equilibrium for Brownian particles. We argue that these results are also relevant for a class of more realistic interactions with both an attractive and a repulsive part, but which decay at infinity.
@@ I.1 Overview
In this paper we consider one-dimensional RTPs interacting via a pairwise polynomial potential of the form $W(r)=-k_{0}\,r^{2}/2+g\,r^{4}/4$, which for $k_{0},g>0$ is repulsive at short distance and attractive at large distance. Although the interaction force $W^{\prime}(r)$ diverges at large distance, which is a priori an unrealistic feature (see however below), it has two advantages from a computational perspective. First, $W^{\prime}(0)=0$, hence it allows for particle crossings and therefore allows to use the extension of the Dean-Kawasaki method Dean ; Kawa to the RTP (while it was shown that this method fails to describe single-file active particles TouzoDBM2023 ). Second, it is simple enough so that the self-consistent equation obeyed by the stationary density in the large $N$ limit, derived in activeRDshort , can be solved analytically. We find that in the steady state the system forms a bound state with a bounded support, which exhibits a transition between a joint support and a disconnected support where particles spontaneously split into two groups. Furthermore, remarkably, we find that the steady state is not unique, and that in some range of parameters it exhibits bistability, as well as a breaking of the parity symmetry. The steady state which is reached at large time is found to depend on some features of the initial conditions. By contrast, both of these features are absent in the case of Brownian particles, for which the Gibbs equilibrium is unique. We note that, although this work focuses on the large $N$ limit of the model, the existence of multiple steady states and the breaking of parity symmetry are already visible for small values of $N$, as we have observed from numerical simulations, which we report in Appendix B.
@@ I.3 Main results
In order to test these analytical predictions, obtained in the limit $N\to+\infty$, we have performed direct numerical simulations of the equation of motion (1). The stationary density is determined by averaging over a large time window. We find a good agreement between theory and numerics already for $N=100$ (see Fig. 3). In particular, both the bistability (see Fig. 5) and the existence of asymmetric steady-states (see Fig. 8) are confirmed by these numerical results. The case of small values of $N$ is discussed in Appendix B.
@@ II.3.1 Relation between $k$ and $k_{0}$
Figure 4: Top left: Plot of $m_{2}$ versus $k$ for different values of $\gamma$. In the limit $\gamma\to 0^{+}$ (black curve), $m_{2}(k)$ has a simple expression given in (47). In this case, the function $m_{2}(k)$ is discontinuous at $k_{c}$, but it is smooth for any $\gamma>0$. Top right: Same plot zoomed around $k_{c}$, and for values of $\gamma$ close to $\gamma_{c}$. While there seems to be a cusp at $k_{c}$ for small $\gamma>0$ when looking at the large scales, the function $m_{2}(k)$ appears to be smooth when zooming sufficiently. Bottom left: Plot of $k$ versus $k_{0}$. When $\gamma$ is smaller than some critical value $\gamma_{c}=0.1787369$..., the function $k_{0}(k)$ becomes non-monotonous close to $k_{c}$. The inverse function $k(k_{0})$ thus becomes multi-valued, leading to the coexistence of 2 stable and one unstable steady states for the same value of $k_{0}$, corresponding to different values of $k$. The value $\gamma_{c}=0.1787369$... was computed numerically, using as a criterion that the minimum of $\frac{dk_{0}}{dk}$ vs $k$ vanishes at $\gamma_{c}$ (it is negative for $\gamma<\gamma_{c}$ and positive for $\gamma>\gamma_{c}$). For $\gamma=0^{+}$, the bistability occurs in the interval $k_{0}\in[6.614585...,9.449407$...]. Bottom right: Same plot zoomed around $k_{c}$, and for values of $\gamma$ close to $\gamma_{c}$. The region where the function $k_{0}(k)$ is non-monotonous seems to be entirely located at $k<k_{c}$.
@@ II.3.3 Bistability
$m_{2}(k)=\frac{1}{2}(y_{1}(k)^{2}+y_{3}(k)^{2})\quad,\quad k\geq k_{c}\;,$ (47)
@@
As a side remark, we note that in the limit $\gamma\to 0^{+}$, the values of $y_{1}$ and $y_{3}$ can be obtained explicitly as a function of the bare interaction parameter $k_{0}$ by inserting the relation (47) into the cubic equation (25) (using that $k=k_{0}-3m_{2}$). For $k<k_{c}$ this gives that $y_{1}$ is the only real root of
@@ III.4 Numerical results
We have tested some of the predictions above in numerical simulations (for $N=100$ particles). One finds that indeed starting from an asymmetric initial condition, e.g. separating the particles into two groups of unequal size at time $t=0$, one can obtain a non-symmetric stationary density, which is well described by the expression (78)-(79) – see Fig. 8 for an example. In addition we have studied numerically small values of $N$ in Appendix B. Interestingly, non-symmetric stationary densities can already be observed for $N=5$. However, they did not arise in our simulations for $N=3$ or $4$, which could be interpreted as a sign that the limiting value $\alpha=1/3$ computed above for $N\to+\infty$ also holds at finite $N$.
@@ IV Conclusion
There are several open problems even for this simple type of model. For instance one would like to be able to solve for the dynamics and obtain the basin of attraction of each steady state. Another question is whether, for finite $N$, the noise can allow for rare transitions between these steady states, and at which rate. One could also try to extend this to higher order degree polynomials (or more general functions) and determine the co-dimension of the stationary manifold.
@@
Acknowledgments. We acknowledge support from ANR Grant No. ANR- 23-CE30-0020-01 EDIPS. We thank Naftali Smith for his interesting comments.
@@ Appendix A Self-consistent equation for the stationary density
$\partial_{t}\rho_{s}=\partial_{x}[(-v_{0}\rho_{d}-\tilde{F}(x,t)\rho_{s}]\;,$ (91)
@@
The equations (91) are identical to the equations of a single RTP, $N=1$, in a effective time dependent force field $\tilde{F}(x,t)$ which depends itself on the time dependent density. One can thus use the known results for this simpler problem, whenever they are available.
@@ Appendix B Double-well interaction for $N=2$, and numerics for small values of NN
These results show that the existence of multiple stationary states is not just an effect of large $N$ but is also present for small values of $N$ (e.g. for $N=5$ one may have either $2$ particles on the left and $3$ on the right or vice versa).